OpenStudy (anonymous):

Which of the following are vertical asymptotes of the function y = 2cot(3x) + 4? Check all that apply. A. x=2pi B. x=pi/3 C. x=+- pi/2 D. x=0

OpenStudy (sidsiddhartha):

\[cotx=cosx/sinx\] so check for whatever \[\sin3x=0\]

OpenStudy (anonymous):

i think domain is all real numbers except x=0 and x= pi/3

OpenStudy (anonymous):

yeah so A and D are definitely the solution

OpenStudy (sidsiddhartha):

look sin3x will be equal to zero for x=pi/3,2pi, and 0 so A,B,D looks good to me

OpenStudy (anonymous):

but yeah if we talk about period, Period = (2pi/3) = (2/3)pi---> pi/2 = x so yes B also gives the solution

OpenStudy (sidsiddhartha):

for x=pi/3 \[\sin(3x)=\sin(3*\frac{ \pi }{ 3 })=\sin \pi=0\] for x=2pi\[\sin(3x)=\sin(2*3\pi)=0\] and for x=0 \[\sin(3*0)=\sin0=0\]

OpenStudy (anonymous):

Thank you!! It was A,B, and D.

OpenStudy (sidsiddhartha):

yup :)